Trace the curve y


2 = (x − 1)(x − 2)(x − 3).


1
Expert's answer
2021-01-05T18:16:58-0500

Solution

"y=(x-1)(x-2)(x-3)-2"

  • As it is easily discoverable,

"(1,-2) ,(2,-2),(3,-2),(0,-8)"

  • For the next calculations it is easy if this is expanded as follows

"y=x^3-6x^2+11x-8"

  • To estimate where would the graph (y) tend as x tends towards +/- infinity, set x to either +/-infinity.

"y\to+\infty" as "x\to+\infty"

"y\to -\infty" as "x \to -\infty"

  • To evaluate the maxima and minima, take the derivative & find the roots

"y'=3x^2-12x+11=0"

  • This gives

"x=2.57" or "x=1.42"

  • Then the corresponding coordinates are

"(2.57,-2.38)" and "(1.42,-1.62)"

  • With all these coordinates, the graph can be traced as follows